# An Idiosyncrasy of Time-discretization in Reinforcement Learning

### v1

@@ 7 Conclusions and Future Work
In this work, we identified and characterized an idiosyncrasy of time-discretization in reinforcement learning. Specifically, a nuance between the definitions of the discrete-time and continuous-time returns when viewing one as a discretization of the other. Our results suggest that when one does not have access to evaluating the integral return via options, one can better align the objectives by shifting the discount factor to begin discounting sooner. This provides unification in that the discrete-time return becomes a relatively straight-forward discretization of the integral return. We strongly emphasize the simplicity of the modification, and how apart from the $\gamma=0$ extreme, such a modification has no loss of generality in discrete-time or with fixed discretization intervals due to proportionality with the conventional discrete-time return. The returns are equivalent as $\gamma^{\Delta}\rightarrow 1$, but as it deviates, it results in better prediction targets in terms integral approximation error, and improves control performance with variable time-discretization. Beyond the integral approximation perspective, the modification has intuitive appeal in that results from catastrophically long delays are attenuated in the return, rather than fully crediting an action for that outcome.
This work was built on an assumption that the reward better aligns with the subsequent time-step, which is often the case in how existing continuous-time environments are set up. The right-point Riemann sum can be viewed as evaluating the integral within the interval for an impulse reward at the subsequent time-step, as would be done by the options framework in a semi-MDP. However, we emphasize that this work still considers problems with arbitrarily dense rewards, but due to discrete sampling, rewards appear as delayed impulse rewards from the agent’s perspective. Should there be additional information about when a reward occurs within an interval, the ideas still generalize in that discounting can be shifted to reflect this information.
Regarding avenues for future work, the integral approximation perspective suggests opportunity to explore return modifications corresponding with other integral approximation techniques. For example, if one were to additionally track predecessor rewards, it opens up the possibility of interpolation-based approximations (e.g., trapezoidal rule). For the case of exponential discounting, we could further leverage a closed-form integral for that term.
@@ Acknowledgements
The authors were generously supported by Amii, NSERC, and CIFAR, and would like to thank Alan Chan and Sungsu Lim for insights and discussions contributing to the results presented in this paper. The authors further thank the reviewers for valuable feedback during the review process.
Appendix A Discrete-time Return Visualization with Variable Interval Sizes
Below we provide a visualization akin to Figure 1, but with variable discretization intervals.
Figure 7: The resulting sum when applying a discrete-time algorithm to a discretized continuous-time domain with variable interval sizes. Note how rectangle heights may fall out of the function’s range within an interval.
Appendix B Servo Reacher Environment Details
@@
where $\Delta_{s}=10^{-4}$ [s] is the simulation discretization granularity, and $A_{t}$ is an input voltage with built-in saturation limits of $\in[-12,12]$ [V]. The output shaft angle is clamped $\in[-1.306,1.306]$ [rad] in accordance with Mahmood et al. (2018b). The physical parameters used are detailed below:

### v2

@@ 7 Conclusions and Future Work
In this work, we identified and characterized an idiosyncrasy of time-discretization in reinforcement learning. Specifically, a nuance between the definitions of the discrete-time and continuous-time returns when viewing one as a discretization of the other. Our results suggest that when one does not have access to evaluating the integral return via options, one can better align the objectives by shifting the discount factor to begin discounting sooner. This provides unification in that the discrete-time return becomes a relatively straight-forward discretization of the integral return. We strongly emphasize the simplicity of the modification and how apart from the $\gamma=0$ extreme, such a modification has no loss of generality in discrete-time or with fixed discretization intervals due to proportionality with the conventional discrete-time return. The returns are equivalent as $\gamma^{\Delta}\rightarrow 1$, but as $\gamma^{\Delta}$ deviates from 1, the right-point return is a better prediction target in terms of integral approximation error and improves control performance with variable time-discretization. Beyond integral approximation, the modification has intuitive appeal in that results from catastrophically long delays are attenuated in the return, rather than fully crediting an action for that outcome.
This work assumed that rewards better align with the subsequent time-step, as is often the case in the setups of existing continuous-time environments. However, should there be domain knowledge suggesting that an environment’s rewards align with some other point in an interval, the ideas generalize in that discounting should be properly exponentiated to reflect this information.
Regarding avenues for future work, the integral approximation perspective suggests opportunity to explore return modifications corresponding to other integral approximation techniques. If one were to additionally track predecessor rewards, it opens up the possibility of interpolation-based approximations like the trapezoidal rule. Notably, Ayoub et al., (2024) concurrently considered trapezoidal approximations of the Monte Carlo return while exploring when to discretize. For the case of exponential discounting, however, we could further leverage that term’s closed-form integral.
@@ Acknowledgements
This research was generously supported by Amii, NSERC, Google Deepmind, and the Pan-Canadian AI Strategy managed by CIFAR. We would like to thank Forte Shinko, Alan Chan, and Sungsu Lim for insights and discussions contributing to the results in this paper, and the reviewers for valuable feedback during the review process.
Appendix A Servo Reacher Environment Details
@@
where $\Delta_{s}=10^{-4}$ [s] is the simulation discretization granularity, and $A_{t}$ is an input voltage with built-in saturation limits of $\in[-12,12]$ [V]. The output shaft angle is clamped $\in[-1.306,1.306]$ [rad] in accordance with Mahmood et al., 2018b . The physical parameters used are detailed below: